You are encouraged to discuss this homework with classmates or with me, but you must write the answers you turn in completely on your own. Keep a copy of your answers to refer to in class. You may turn in answers on paper or in a digital copy (in a single file) submitted at Blackboard. If you turn in a paper copy by the deadline, it will be graded as your final copy. All answers must be clearly legible.

This assignment refers to the following quotes from “Climate Change Isn’t the End of the World” by Henderson and Cochrane. The whole article from the Wall Street Journal 7/31/17 is at a link next to the link with this assignment.

A. “To arrive at a wise policy response, we first need to consider how much economic damage climate change will do. Current models struggle to come up with economic costs commensurate with apocalyptic political rhetoric. Typical costs are well below 10% of gross domestic product in the year 2100 and beyond. That’s a lot of money—but it’s a lot of years, too. Even 10% less GDP in 100 years corresponds to 0.1 percentage point less annual GDP growth. Climate change therefore does not justify policies that cost more than 0.1 percentage point of growth. If the goal is 10% more GDP in 100 years, pro-growth tax, regulatory and entitlement reforms would be far more effective.”

B. “Climate change need not endanger anyone. If it did—and you do hear such claims—then living in hot Arizona rather than cool Maine, or living with Louisiana’s frequent floods, would be considered a health catastrophe today.”

C. “Climate policy must compete with other long-term threats for always-scarce resources. Facing this reality, some advocate that we buy some “insurance.” Sure, they argue, the pro- jected economic cost seems small, but it could turn out to be a lot worse. But the same argument applies to any possible risk. If you buy overpriced insurance against every potential danger, you soon run out of money. You can sensibly insure only when the premium is in line with the risk—which brings us back where we started, to the need for quantifying probabilities, costs, benefits and alternatives. And uncertainty goes both ways. Nobody forecast fracking, or that it would make the U.S. the world’s carbon-reduction leader. Strategic waiting is a rational response to a slow-moving uncertain peril with fast-changing technology.”

Problems:

1.[5] Assume that US GDP would grow at 2% per year over then next 100 years if the quantity of atmospheric greenhouse gasses never rose above its current level. Use the mathematics of a constant growth rate to show how the authors find in quote A that 10% lower GDP in 100 years corresponds approximately to 0.1 percentage points lower annual GDP growth between now and then (in other words, to 1.9% growth per year instead of 2%). Explain all the steps of your reasoning.
Answer #1

Problem : How a 0.1% lower annual GDP growth between now and the next hundred years leads to a 10% lower GDP in next hundred years.

Method of solving : Using constant GDP growth rate formula .

Solution :

Present growth % = (1+gr )n

Where ,

gr = growth rate

n = Number of years

In the given question ,

n = 100

1) present growth percentage when the annual GDP growth rate is 2%

Growth rate = (1+0.02)100=7.24%

2) present growth percentage when the annual GDP growth rate is 1.9 %

Growth rate = (1+0.019)100=6.57%

Therefore the change in present growth rate after hundred years after comparison of both the cases will be:-

Change in annual growth rate = (7.24-6.57)/6.57*100

=10.19%

Observations :-

A mere change of 0.1% reduction in in the annual growth rate would compound to a a 10% in the coming hundred years